OpenStudy (anonymous):

a. 1+2+3+4+....+998 b. 1+3+5+7+...+1003 a. The sum of the sequence is? b. The sum of the sequence is?

OpenStudy (anonymous):

a.=Sum from n=1 to 998 [n]=(998+1)*998/2=498501 b.= Sum from n=1 to 52 [2n-1]=2*(Sum from n=1 to 52)-52= 2*((52+1)*52/2)-52=(53)*52-52=52*52=2704

OpenStudy (anonymous):

Wasn't it 1 3 5 ... 103? My bad :s b.= Sum from n=1 to 502 [2n-1]=2*(Sum from n=1 to 502)-502= 2*((502+1)*502/2)-502=(503)*502-502=502*502=252004